# GraspLDM: Generative 6-DoF Grasp Synthesis using Latent Diffusion Models

### v1

@@ V Conclusions
Learning generalizable representations for grasp generation is fundamental to robotic grasping. In this work, we propose GraspLDM- a novel generative framework for object-centric 6-DoF grasp synthesis using latent diffusion. Our architecture allows de-noising diffusion models to be used as expressive priors in the latent space of VAE s. This enables efficient learning of the complex distribution of object-centric 6-DoF grasp poses on point clouds. We conduct large-scale simulation tests to show that GraspLDM outperforms baseline methods and provides a 78% median success rate on our test set of 400 objects from 63 categories. GraspLDM shows good grasp generation performance while scaling favorably to large object sets compared to existing methods using generative modeling. Further, we validate our pipeline by training GraspLDM models on single-view point clouds generated in simulation and testing them in a real-world setup. GraspLDM demonstrates favorable transfer to the real world and provides 80% success rate on 16 test objects from single-view depth images. We further demonstrate the flexibility of our architecture to train efficient task-specific models and to use fast reverse diffusion samplers for downstream applications. GraspLDM can be extended to work with complex task conditionings and cluttered scenes.
Figure 7: Test object set containing objects with diverse visual, geometric and surface properties for real-world evaluation (left) and example grasps executed by GraspLDM (right)
Acknowledgments
This work is supported by the Fonds National de la Recherche (FNR) Industrial Fellowship grant (15799985) and Redwire Space Europe.

### v2

@@ VI Conclusions
Learning generalizable representations is fundamental to robotic grasping in the real world. In this work, we propose GraspLDM —a novel generative framework to learn object-centric representations for 6-DoF grasp synthesis. Our architecture allows denoising diffusion models to be used as expressive priors in the latent space of VAE s. This enables efficient learning of the complex distributions of 6-DoF grasp poses on object-centric point clouds. We conduct large-scale simulation tests to show that GraspLDM outperforms baseline methods and provides a 78% median success rate on our test set of 400 objects from 63 categories. GraspLDM demonstrates better grasp generation performance while scaling favorably to large object sets compared to existing generative models. We also demonstrate GraspLDM’s ability to efficiently train task-specific models and use fast reverse diffusion samplers for downstream applications. Further, we validate our pipeline by training GraspLDM models on single-view point clouds generated in simulation and testing them on two different real-world setups. GraspLDM demonstrates effective sim-to-real transfer, achieving 78% and 80% success rates across 80 grasp attempts on 16 diverse objects in two different real-world robotic setups. GraspLDM outperforms contemporary generative models and matches the performance of state-of-the-art non-generative models while offering greater flexibility.